Shift summary for day shift, 8/Jan/00


Subject: Shift summary for day shift, 8/Jan/00
From: Vivian O'Dell (odell@fnal.gov)
Date: Fri Feb 04 2000 - 15:10:43 CST


               Shift summary for day shift, 8/Jan/00

On Shift: PG, AA

Hours of beam available at all intensities: 8.3h
Hours of beam available at nominal intensity: 8.0h
Hours of E799 data (special runs not included): 6.5h
Hours of special runs (laser scans, muon runs...): 0.0h

Accounting of lost time:

A tape drive failure and subsequent daq problems to start a run, continued
from the end of the owl shift. Resetting pipeline, ktevana errors, trigger
busy, many many KFREND's.
========================================================================

Runs ended this shift:

15441 BEAM_799 ongoing

Tapes brought to FCC this shift:
  (including tapes for on-line split)

UPA301-306, UPB301-306, UPC301-306, UPS101-102,
UPF073, UPK039, UPH054, UPM054

========================================================================

Accesses during shift:

none.

========================================================================

Overview of shift:

Began with a lot of difficulty starting a new run for more than an hour.
Relatively smooth 799 data-taking after that, occasional trigger errors.

========================================================================

Description of Shift:

07:00-8:10 Owl shift still struggling to start a new run. Several
            drt_fresh_start's failed. We discovered tape drive3 in ktev2
            was in error. After removing the stuck tape (by four of us!),
            replacing with a new tape, cycling the power on ktev2 drives,
            drt_fresh_start and initialization, we failed to start the run
            again. Finally, Ed tracked the problem down to the fast bus
            creates which were not communicating with the daq. Reset all
            the crates and finally started a new run after more than an
            hour. AG who has been helping along is dismissed!

Run Spill Time Description

             7:40 onsplit tape UPH055 requested. This replaces the
                       one which was already on ktev2 drive5 and got
                       screwed when we cycled the power.
 
15441 1 8:10 Run starts.

       43 9:08 KTEVANA error, KFREND.

       64 9:35 CsI PED problems and laser rms too wide error.
                       Channel 2085 is highly inefficient. Resetting the
                       pipeline resolved it. Spills 64-70 are affected.

       84 10:05 No beam for spills 78-84.

      120 10:50 Trigger error, KFREND.

      139 11:15 No beam for spills 132-139.

            11:45 Onsplit tape UPM055 mounted on ktev1.

            12:45 Trigger hanging. Cycled stream 5 repeater
                       downstairs and upstairs and several KFRENDs.
                       Affected spills 195-205.

      224 13:18 trigger busy again. 3 KFREND's.

            13:40 Trigger hanging. Cycled stream 5 repeater
                       downstairs several KFREND. Affected spills 236-243.

            14:30 noticed the life time is dropping and chambers
                       drawing too much current. Call the control room to
                       correct the spill structure.

            15:00 MCR corrected beam. Live time improving.
                   
========================================================================

Follow up Notes on Previous Shift

========================================================================

Items to Follow up on Next Shift

Rick K. would like to have BA channel 141 updated. Instructions in the
logbook page 210.

From Leo: When tapes full, swap the polarity of both magnets, take a muon
run with BA high voltage as per Mike Arenton. Take hadrons after that.



This archive was generated by hypermail 2b27 : Fri Feb 04 2000 - 15:10:45 CST